Prove that the equation normal to the rectangular hyperbola xy = c^2 at the point P(ct, c/t) is t^3*x - ty = c(t^4 - 1)
1
Expert's answer
2021-02-02T05:21:25-0500

The equation of the normal to the graph of the function "y = y(x)" at a point with abscissa "x_0" is the following:


"y-y(x_0)=-\frac{1}{y'(x_0)}(x-x_0)"


In our case, "y=\frac{c^2}{x}, \ \ x_0=ct,\ y(x_0)=\frac{c}{t}". It follows that "y'=-\frac{c^2}{x^2}". Therefore, the equation of the normal at point "x_0=ct" is


"y-\frac{c}{t}=-\frac{1}{-\frac{c^2}{c^2t^2}}(x-ct)" which is equivalent to


"y-\frac{c}{t}=t^2(x-ct)" and to


"ty-c=t^3x-ct^4".


Finally, we have the equation of the normal in the following form:


"t^3x-ty=c(t^4-1)"